OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took one-third the number of tickets, that her cousin took, which was half the number Heather took. If her brother took 8 tickets, how many tickets did Heather take? 8 24 48 72

OpenStudy (theeric):

Hi! So, I'd make variables for each person. How does that sound?

OpenStudy (anonymous):

@theEric I got 8 is that right?

OpenStudy (theeric):

That's not what I got, so I don't think so.. How did you work it out?

OpenStudy (anonymous):

@theEric someone else tried helping and told me he did this.. A=Brother ->8 tickets B=cousin C=Heather 1/3B=A B=C/2

OpenStudy (theeric):

I agree with that.

OpenStudy (anonymous):

@theEric then its 8 right? if not what is he trying to say in the epression??

OpenStudy (anonymous):

is 48

OpenStudy (theeric):

So, we know A and we want to know C. So we substitute C/2 in for the B in your top equation, 1/3 * B=A. We get 1/3 *C / 2 = A. That is, C/6=A, and C=6A, right?

OpenStudy (theeric):

@francoj22 http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@francoj22 @theEric Thank you both!

OpenStudy (anonymous):

@theEric no he didnt tell me the answer he explained it and i asked him dont blame it on him

OpenStudy (theeric):

You're welcome! You did the work too! :) You did the tough word-problem part, and maybe just made a small algebra mistake :) Okay, I got ya! Thanks for letting me know!

OpenStudy (anonymous):

@theeric thanks for helping

OpenStudy (theeric):

No problem! Take care!